OpenStudy (anonymous):

Factor the polynomial function f(x). Then solve the equation f(x)= 0. f(x) x^4 -x^3 -19x^2 +49x -30

OpenStudy (campbell_st):

I'd say try and find 4 factors of -30 use the factor theorem and try \[\pm 1, \pm 2, \pm 3 \pm 5\] with 4 factors you'll need a odd number of negative factors so start by finding P(1), then P(2), P(3) and P(5) if any equal to zero... then you have a zero... and can write the factor...

hero (hero):

\[f(x) = x^4 - x^3 - 19x^2 + 49x - 30\] \[=x^3(x - 1) - x(19x - 49) - 30\] \[=x^3(x - 1) - x(x - 1 + 18x - 48) - 30\] \[=x^3(x - 1) - x(x - 1) - 18x^2 + 48x - 30\] \[=x^3(x - 1) - x(x - 1) - 18x^2 + 48x - 30\] \[=(x - 1)(x^3 - x) - 18x^2 + 48x - 30\] \[=x(x - 1)(x^2 - 1) - 6(3x^2 - 8x + 5)\] \[=x(x - 1)(x + 1)(x - 1) - 6(x - 1)(3x - 5)\] \[=(x - 1)(x(x + 1)(x - 1) - 6(3x - 5))\] \[=(x - 1)(x(x + 1)(x - 1) - 6(x - 1 + 2x - 4))\] \[=(x - 1)(x(x + 1)(x - 1) - 6(x - 1) - 12x + 24)\] \[=(x - 1)(x(x + 1)(x - 1) - 6(x - 1) - 12(x - 2))\] \[=(x - 1)((x - 1)(x(x + 1) - 6) - 12(x - 2))\] \[=(x - 1)((x - 1)(x^2 + x - 6) - 12(x - 2))\] \[=(x - 1)((x - 1)(x + 3)(x - 2) - 12(x - 2))\] \[=(x - 1)((x - 2)(x - 1)(x + 3) - 12))\] \[=(x - 1)(x - 2)(x^2 + 2x - 3 - 12)\] \[=(x - 1)(x - 2)(x^2 + 2x - 15)\] \[0=(x - 1)(x - 2)(x - 3)(x + 5)\] x - 1 = 0 x - 2 = 0 x - 3 = 0 x + 5 = 0